# The statistical thermodynamics of generative diffusion models: Phase transitions, Symmetry breaking and critical instability

### v2

@@ 7 Associative memory and Hopfield networks
Associative memory networks are energy-based learning systems that can store patterns (i.e. memories) as meta-stable states of a parameterized energy function (Hopfield,, 1982; Abu-Mostafa and Jacques,, 1985; Krotov,, 2023). There is a substantial body of literature on the thermodynamic properties of associative memory networks (Strandburg et al.,, 1992; Volk,, 1998; Marullo and Agliari,, 2020). The original associative memory networks, also known as Hopfield networks, are defined by the energy function $E(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}W\boldsymbol{x}$ under the constraints of binary entries for the state vector. In a Hopfield network, a finite number of training patterns $\boldsymbol{y}_{j}$ are encoded into a weight matrix $W=\sum_{j}\boldsymbol{y}_{j}\boldsymbol{y}_{j}^{T}$, which usually gives the correct minima when the number of patterns is on the order of the dimensionality. Associative memory networks can reach much higher capacity by using exponential energy function (Krotov and Hopfield,, 2016; Demircigil et al.,, 2017; Krotov,, 2023). For example, (Ramsauer et al.,, 2021) introduces the use of the following function
@@ 8.1 Memorization as ’condensation’
From our theoretical framework, we can find this ’critical time’ $t_{c}$ by studying the random partition function given in Eq. 34 in the thermodynamic limit. This can be done by evaluating the ’participation ration’
@@
For high values of $d$, the distribution of $\boldsymbol{x}\cdot\boldsymbol{y}_{j}$ is approximately a centered Gaussian with variance ${\left\lVert x\right\rVert}_{2}^{2}\nu^{2}M/2$ and, which $\phi(\boldsymbol{y})$ is spherically symmetric, it does not depend on the direction of $\boldsymbol{x}$. We can therefore re-write the expression as a standard REM model:
@@
where $\tilde{E}_{j}\sim\mathcal{N}(0,M/2)$ and $\tilde{\beta}=\nu\beta(t){\left\lVert\boldsymbol{x}\right\rVert}_{2}$. For this model, we know that the critical condensation temperature is at $\beta_{c}=2\sqrt{\log{2}}$ (see Mézard et al., (1987)), which leads to the critical time
$t_{c}(\boldsymbol{x})=\frac{\nu{\left\lVert x\right\rVert}_{2}}{2\sigma^{2}\sqrt{\log{2}}}~{}.$ (38)
@@
$\mathbb{E}\!\left[Y(\boldsymbol{x},t)\right]=\left\{\begin{array}[]{lr}0,&\text{if }\beta(t)\leq\beta_{c}\\ 1-\frac{\beta_{c}}{\beta(t)},&\text{if }\beta(t)>\beta_{c}\end{array}~{}.\right\}$ (39)

### v4

@@ 9 Associative memory and Hopfield networks
We will now move back to the standard mean-field formulation of generative diffusion and discuss its connection with associative memory networks. Associative memory networks are energy-based learning systems that can store patterns (i.e. memories) as meta-stable states of a parameterized energy function (Hopfield,, 1982; Abu-Mostafa and Jacques,, 1985; Krotov,, 2023). There is a substantial body of literature on the thermodynamic properties of associative memory networks (Strandburg et al.,, 1992; Volk,, 1998; Marullo and Agliari,, 2020). The original associative memory networks, also known as Hopfield networks, are defined by the energy function $E(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}W\boldsymbol{x}$ under the constraints of binary entries for the state vector. In a Hopfield network, a finite number of training patterns $\boldsymbol{y}_{j}$ are encoded into a weight matrix $W=\sum_{j}\boldsymbol{y}_{j}\boldsymbol{y}_{j}^{T}$, which usually gives the correct minima when the number of patterns is on the order of the dimensionality. Associative memory networks can reach much higher capacity by using exponential energy function (Krotov and Hopfield,, 2016; Demircigil et al.,, 2017; Krotov,, 2023). For example, (Ramsauer et al.,, 2021) introduces the use of the following function
@@ 10.1 Memorization as ’condensation’
From our theoretical framework, we can find this ’critical time’ $t_{\text{cond}}$ by studying the random partition function given in Eq. 43 in the thermodynamic limit. This can be done by evaluating the ’participation ration’
@@
For high values of $d$, the distribution of $\boldsymbol{x}\cdot\boldsymbol{y}_{j}$ is approximately a centered Gaussian with variance ${\left\lVert x\right\rVert}_{2}^{2}\nu^{2}M/2$ and, which $\phi(\boldsymbol{y})$ is spherically symmetric, it does not depend on the direction of $\boldsymbol{x}$. We can therefore approximately re-write the expression as a standard REM model:
@@
where $\tilde{E}_{j}\sim\mathcal{N}(0,M/2)$ and $\tilde{\beta}=\nu\beta(t){\left\lVert\boldsymbol{x}\right\rVert}_{2}$. Note that this formulation is not exact at the thermodynamic limit as it ignores the residual effect of non-Gaussianity in the distribution. A more rigourous analysis can be found in (Lucibello and Mézard,, 2024) (Appendix C) in the context of modern Hopfield networks.
For the standard REM model, we know that the critical condensation temperature is at $\beta_{c}=2\sqrt{\log{2}}$ (see Mézard et al., (1987)), which leads to the critical time
$t_{\text{cond}}(\boldsymbol{x})=\frac{\nu{\left\lVert x\right\rVert}_{2}}{2\sigma^{2}\sqrt{\log{2}}}~{}.$ (47)
@@
$\mathbb{E}\!\left[Y(\boldsymbol{x},t)\right]=\left\{\begin{array}[]{lr}0,&\text{if }\beta(t)\leq\beta_{c}\\ 1-\frac{\beta_{c}}{\beta(t)},&\text{if }\beta(t)>\beta_{c}\end{array}\right\}~{}.$ (48)